OpenStudy (anonymous):

f(x) = x^2 -2x +3 find f(-x)

OpenStudy (anonymous):

\(f(-x)=(-x)^2-2\times(-x)+3=x^2+2x+3\)

OpenStudy (anonymous):

so i just multiply the -x with all those who contain an x? since youve excluded the 3

OpenStudy (anonymous):

DO NOT MULTIPLY .JUST REPLACE X WITH -X

OpenStudy (anonymous):

replace the xs with a -x, except for three... lol i feel stupid.. :p

OpenStudy (anonymous):

wait.. why did i just replace the second x and not the first?

OpenStudy (anonymous):

OH becuase its squared,..

OpenStudy (anonymous):

yes because \((-x)^2=(-x)(-x)=x^2\)

OpenStudy (anonymous):

oh yeahhhh.. wow, i feel stupid again :o

OpenStudy (anonymous):

i think i get it, i'l go marinate this thing for a few more mins. thanks a bunch both of yall! :p